### v1

# TIPS Over Tricks: Simple Prompts for Effective Zero-Shot Anomaly Detection

Abstract
Anomaly detection identifies departures from expected behavior in safety-critical settings. When target-domain normal data are unavailable, zero-shot anomaly detection (ZSAD) leverages vision–language models (VLMs). However, CLIP’s coarse image–text alignment limits both localization and detection due to (i) spatial misalignment and (ii) weak sensitivity to fine-grained anomalies; prior work compensates with complex auxiliary modules yet largely overlooks the choice of backbone. We revisit the backbone and use TIPS—a VLM trained with spatially aware objectives. While TIPS alleviates CLIP’s issues, it exposes a distributional gap between global and local features. We address this with decoupled prompts—fixed for image-level detection and learnable for pixel-level localization—and by injecting local evidence into the global score. Without CLIP-specific tricks, our TIPS-based pipeline improves image-level performance by 1.1–3.9% and pixel-level by 1.5–6.9% across seven industrial datasets, delivering strong generalization with a lean architecture. Code is available at github.com/AlirezaSalehy/Tipsomaly.
Index Terms—  Zero-shot Anomaly Detection, Vision-language Model
1 Introduction
Anomaly detection seeks rare deviations in visual data and is pivotal in safety-critical settings such as industrial inspection and medical imaging [27, 7]. Conventional approaches commonly assume access to sufficient in-domain normal data and often model it via deviations in feature distribution [28, 25, 24, 23] or reconstruction [22, 12]. However, when labeled in-domain normal samples are unavailable—due to privacy or scarcity—zero-shot anomaly detection (ZSAD) becomes essential. Recent vision–language models (VLMs) such as CLIP [21] make this setting practical by coupling text prompts with robust transferable visual representations, enabling anomaly detection in unseen domains without target-domain training data [13, 9].
However, CLIP’s pretraining poses a bottleneck for ZSAD. Its contrastive objective does not enforce patch-level alignment—neither among patch embeddings nor between patches and text [21]. This weak spatial coherence and patch–text grounding lead to degraded localization of fine-grained cues [17], which in turn hampers both detection and pixel-level performance [6].
To mitigate these issues, prior ZSAD methods introduce non-learnable refinements (“VV” attention [34, 6]) and learnable modifications to the vision encoder (trainable visual prompts [4], feature adapters [5]). Although helpful, these adaptations cannot fully restore the spatial alignment and patch–text correspondence absent in CLIP inherently, and increase architectural complexity. Moreover, previous analyzes suggest that learnable adaptations can overfit train data, weakening the cross-domain generalization required for zero-shot anomaly detection [26]. For example, AdaCLIP’s [4] joint tuning of text and vision encoders can improve image-level detection, but transfers poorly to pixel-level localization, underscoring the limits of CLIP-centric adaptations.
In this paper, we revisit an underexplored lever for ZSAD—the backbone—and investigate TIPS’s [19] capabilities, a more spatially aware model for the task, showing it sidesteps several CLIP limitations and enables a simpler pipeline without complex adaptations. However, TIPS is trained with a generic vision–language objective, making direct utilization suboptimal for anomaly detection. We evaluated learnable prompts and fixed textual templates and found a consistent distributional gap between local and global features: when learnable prompts are trained with an additional global objective (normal vs. anomalous classification) alongside the local loss, image-level performance improves but pixel-level alignment degrades, indicating a mismatch between local and global anomaly representations.
To address this, we adopt decoupled prompting: fixed textual prompts for image-level scoring, and learnable prompts optimized only with a local loss for pixel-level segmentation. We also leverage TIPS’s two global embeddings—the spatial token and the object-centric token—and assess their ability to discriminate anomalies and aggregate local cues; we then complement the global score with local evidence. Our framework Tipsomaly improves state-of-the-art image-level performance by 1.1–3.9% and pixel-level metrics by 1.5–6.9% across 14 industrial and medical datasets.
Fig. 1: Framework of Tipsomaly for zero-shot anomaly detection. An input image is encoded by the TIPS vision encoder $\Phi_{v}$ into dense patch features $Z^{\mathrm{M}}$ and two global tokens, $g_{o}^{i}$ (object-centric) and $g_{s}^{i}$ (spatial). We use decoupled prompting: the fixed detection prompt set $P_{f}$ yields fixed text prototypes $G_{f}$ for image-level decisions, while the learnable localization prompt set $P_{l}$ produces $G_{l}$ for pixel-level localization. The anomaly map $\hat{S}$ is computed via patch–text similarities between $Z^{\mathrm{M}}$ and $G_{l}$, then upsampled and smoothed to image resolution. For image-level detection, spatial token $g_{s}^{i}$ is scored against $G_{f}$, then the result is added to the strongest local evidence, $\max(\hat{S}_{a})$, to obtain anomaly score $\hat{y}$.
2 Method
2.1 Problem Definition
In anomaly detection, a dataset is composed of triplets $\mathcal{D}=\{(x,y,S)\}$, where each image $x\in\mathbb{R}^{C\times H\times W}$ is paired with an image-level label $y\in\{0,1\}$ and a pixel-level mask $S\in\{0,1\}^{H\times W}$, with $1$ indicating abnormality. Let $M_{\theta}$ denote a pretrained vision–language model (e.g., CLIP, TIPS) with parameters $\theta$. In the zero-shot setting, the goal is to detect and localize anomalies in an unseen domain $\mathcal{D}_{\mathrm{unseen}}$ without using any target-domain samples during optimization. To this end, we may adapt the model using a source dataset $\mathcal{D}_{\mathrm{src}}$ drawn from an auxiliary domain disjoint from $\mathcal{D}_{\mathrm{unseen}}$ by learning prompts $P$. At test time, for each $x\in\mathcal{D}_{\mathrm{unseen}}$ the adapted model produces an image-level anomaly score $\hat{y}(x)\in[0,1]$ and a pixel-level anomaly map
$\hat{S}(x)\in[0,1]^{H\times W}$:
$(\hat{y}(x),\,\hat{S}(x))\;=\;M_{\theta}(x,P).$
thereafter the final Image- and pixel-level decisions are obtained by thresholding these scores.
2.2 Feature Extraction
To produce pixel- and image-level anomaly scores for an image $x\in\mathbb{R}^{C\times H\times W}$, we extract visual and textual features. As shown in Fig. 1, TIPS—similar to CLIP—comprises a vision encoder $\Phi_{v}$ and a text encoder $\Phi_{t}$. The vision encoder outputs two global tokens—object-centric $g_{o}^{i}\in\mathbb{R}^{D}$ and spatial $g_{s}^{i}\in\mathbb{R}^{D}$—used for image-level decisions, along with dense patch embeddings $Z^{\mathrm{M}}\in\mathbb{R}^{N\times D}$ for pixel-level localization. The text encoder produces embeddings for normal and abnormal states from two prompt sets: fixed detection prompts $P_{f}$ aiming for image-level detection and learnable localization prompts $P_{l}$ specializing the text representations for fine-grained localization. This decoupled prompting addresses a distributional gap between the global tokens and the patch features of the vision encoder observed in our study.
2.3 Fixed Detection Prompts
We compose each fixed prompt from three parts: a template preamble, a state phrase, and the class token. The preamble is drawn from a compact set of generic inspection templates (e.g., a cropped photo of a {}), and the state phrase specifies normal vs. abnormal semantics (e.g., flawless {} vs. damaged {}). Given a class name $c$, we instantiate prompts by the Cartesian product
$P_{f}\;=\;\{\,t(s(c))\;\mid\;t\in\mathcal{T},\ s\in\mathcal{S}\,\},$
and split them into normal and abnormal subsets $P_{f}^{\mathrm{n}}$ and $P_{f}^{\mathrm{a}}$. Passing each prompt $p\in P_{f}$ through the text encoder $T(\cdot)$ yields an embedding $T(p)\in\mathbb{R}^{D}$. We then average embeddings within each subset to obtain fixed text prototypes and denote $G_{f}=\{g_{f}^{\mathrm{n}},\,g_{f}^{\mathrm{a}}\}$. These fixed prototypes are used for image-level detection.
2.4 Learnable Localization Prompts
We observed that fixed prototypes offer strong image-level alignment but fail to encode local cues needed for patch-level alignment. Hence, for localization we learn two class-agnostic prompt token sets on $\mathcal{D}_{\mathrm{src}}$, $T^{\mathrm{n}}=\big[t^{\mathrm{n}}_{1},\dots,t^{\mathrm{n}}_{E}\big]$ and $T^{\mathrm{a}}=\big[t^{\mathrm{a}}_{1},\dots,t^{\mathrm{a}}_{E}\big]$, with $t^{(\cdot)}_{(\cdot)}\in\mathbb{R}^{D}$. We denote the resulting localization prompt set by
$P_{l}\;=\;\big\{\,[T^{\mathrm{n}},~\text{``object''}],\;[T^{\mathrm{a}},~\text{``damaged''},~\text{``object''}]\,\big\}.$
Passing each element of $P_{l}$ through the text encoder $T(\cdot)$ yields the normal/abnormal localization prototypes $G_{l}=\{g^{n}_{l},\,g^{a}_{l}\}$ which used to score patch embeddings $Z^{\mathrm{M}}$ via similarity.
2.5 Training Objective & Inference
Having obtained the textual embeddings specialized for detection $G_{f}$ and localization $G_{l}$ and visual embeddings $e\in\{g_{o}^{i},g_{s}^{i},Z_{j,k}\}$, where $Z_{j,k}$ denotes $(j_{th},k_{th})$ patch embedding of unflattened local feature $Z^{M}$. for input image $x$, we can now compute the likelihood of each visual embedding belonging to the anomalous class ($p_{a}$) by applying the Softmax function to similarity scores:
$p\!\left(e,G_{(\cdot)}\right)=\operatorname{softmax}\!\left(\frac{G_{(\cdot)}^{\top}e}{\tau}\right),\quad p_{a}=\bigl[p(e,G_{(\cdot)})\bigr]_{a}$ (1)
where the temperature $\tau$ is approximately 0.0042 according to TIPS details [19]. We denote the probability of a visual embedding $e$ being abnormal, $p_{a}(e,G)$, as its anomaly score.
Train. We learn the normal/abnormal patch-level visual distribution via learnable prompts $P_{l}$, while keeping the vision encoder $\Phi_{v}$ and text encoder $\Phi_{t}$ frozen and employing local loss function, Focal [18] and Dice [29] loss. For output feature map $Z^{M}$, we compute normal and anomaly maps $\hat{S}_{n}$ and $\hat{S}_{a}$ based on the equation 1, then apply bilinear upsampling $\text{Up}(\cdot)$ to match the anomaly mask $S\in\mathbb{R}^{H\times W}$:
$L=\text{Focal}\left(\text{Up}([\hat{S}_{n},\hat{S}_{a}]),S\right)+\text{Dice}\left(\text{Up}([\hat{S}_{n},\hat{S}_{a}]),S\right).$
Inference. For each input image $x$, after computing visual outputs including local features $Z^{M}$ and global embeddings $g_{o}^{i}$ and $g_{s}^{i}$ alongside detection and localization textual embeddings $G_{f}$ and $P_{l}$ in forward pass, low-resolution anomaly maps $\hat{S}_{a}$ is calculated then bilinearly upsampled and smoothed with Gaussian filter to obtain final anomaly map $\hat{S}$:
$\hat{S}_{a,(j,k)}=p_{a}(Z^{M}_{j,k},G_{l}),\quad\hat{S}=Gaus(Up(\hat{S}_{a})).$
where $\hat{S}_{a,(j,k)}$ is anomaly score for each patch at location $(j,k)$.
To calculate anomaly score $\hat{y}$, we first compute the image-level score $p_{a}(g_{s}^{i},G_{f})$ by comparing the global embedding $g_{s}^{i}$ with the abnormal/normal global prototypes $g^{a}_{f}$ and $g^{n}_{f}$ (collected in $G_{f}$). We then inject strongest local evidence via the maximum pixel-level anomaly score $\max(\hat{S}_{a})$. The final score is
$\hat{y}=p_{a}(g_{s}^{i},G_{f})+\max(\hat{S}_{a}).$
3 Experiments
We evaluate on diverse 14 real-world anomaly detection datasets covering industrial defects and medical abnormalities, including MVTec-AD [2], VisA [35], DTD-Synthetic [1], SDD [30], BTAD [20], DAGM [33], and MPDD [14] for industrial and ISIC [10], CVC-ColonDB [31], CVC-ClinicDB [3], TN3K [8], BrainMRI [15], HeadCT [16], and BR35H [11] for medical domain. Following prior works [34, 4], we train on MVTec and test generalization on the others; for MVTec evaluation, we use VisA. In both cases, the categories are distinct and non-overlapping with those in the other datasets. For metrics, we follow previous studies [5, 34], using AUROC, AP, and F1-max for image-level and AUROC, AUPRO, and F1-max for pixel-level, and report the mean over categories as the dataset-level score.
Table 1: Comparison of ZSAD methods across industrial and medical domains. The best performance is highlighted in bold, while the second-best is underlined. Our method consistently outperforms baselines across different domains, showing strong zero-shot generalization.
Image-level $\uparrow$ (AUROC, AP, F1-max)
Dataset VAND [5] AnomalyCLIP [34] AdaCLIP [4] Ours
MVTec (86.1, 93.5, 88.9) (91.5, 96.2, 92.7) (89.2, 95.7, 90.6) (93.4, 96.1, 92.9)
VisA (78.0, 81.4, 80.7) (82.1, 85.4, 80.4) (85.8, 79.0, 83.1) (87.7, 90.9, 84.8)
MPDD (73.0, 80.2, 76.0) (77.0, 82.0, 80.4) (76.0, 80.2, 82.5) (74.7, 76.8, 79.7)
BTAD (73.6, 68.6, 82.0) (88.3, 87.3, 83.8) (88.6, 93.8, 88.2) (95.0, 95.5, 91.4)
KSDD (79.8, 71.4, 85.2) (84.7, 80.0, 82.7) (97.1, 89.6, 90.7) (97.8, 94.7, 90.1)
DAGM (94.4, 83.8, 91.8) (97.5, 92.3, 90.1) (99.1, 88.5, 97.5) (99.7, 98.1, 96.9)
DTD (94.6, 95.0, 96.8) (93.5, 97.0, 93.6) (95.5, 97.3, 94.7) (99.4, 99.7, 99.1)
Average (81.6, 82.0, 85.9) (87.8, 88.6, 87.2) (90.2, 89.2, 89.6) (92.5, 93.1, 90.7)
HeadCT (89.2, 89.5, 82.1) (93.4, 91.6, 90.8) (91.8, 90.6, 84.1) (92.7, 89.4, 87.4)
BrainMRI (89.6, 91.0, 88.5) (90.3, 92.2, 90.2) (93.5, 95.6, 89.7) (92.4, 95.2, 88.3)
Br35H (91.4, 91.9, 84.2) (94.6, 94.7, 89.1) (92.3, 93.2, 85.3) (93.2, 93.6, 85.6)
Average (90.1, 90.8, 84.9) (92.8, 92.8, 90.0) (92.5, 93.1, 86.4) (92.8, 92.7, 87.1)
Pixel-level $\uparrow$ (AUROC, AUPRO, F1-max)
Dataset VAND [5] AnomalyCLIP [34] AdaCLIP [4] Ours
MVTec (87.6, 44.0, 39.8) (91.1, 81.4, 39.1) (88.7, 37.8, 43.4) (90.9, 84.0, 43.8)
VisA (94.2, 86.8, 32.3) (95.5, 87.0, 28.3) (95.5, 72.9, 37.7) (95.9, 88.2, 31.5)
MPDD (94.1, 83.2, 30.6) (96.5, 88.7, 34.2) (96.1, 62.8, 34.9) (95.9, 86.5, 36.1)
BTAD (60.8, 25.0, 38.4) (94.2, 74.8, 49.7) (92.1, 20.3, 51.7) (96.7, 84.7, 56.2)
KSDD (79.8, 65.1, 56.2) (90.6, 67.8, 51.3) (97.7, 33.8, 54.5) (99.5, 97.9, 58.4)
DAGM (82.4, 66.2, 57.9) (95.6, 91.0, 58.9) (91.5, 50.6, 57.5) (97.2, 94.3, 65.3)
DTD (95.3, 86.9, 72.7) (97.9, 92.3, 62.2) (97.9, 72.9, 71.6) (99.3, 96.0, 70.3)
Average (84.9, 65.3, 46.8) (94.5, 83.3, 46.2) (94.2, 50.1, 50.2) (96.5, 90.2, 51.7)
ISIC (89.5, 77.8, 71.5) (89.7, 78.4, 70.6) (90.3, 54.7, 72.6) (94.0, 87.0, 80.2)
ColonDB (78.4, 64.6, 29.7) (81.9, 71.3, 37.3) (82.6, 66.0, 36.1) (84.6, 75.0, 41.3)
ClinicDB (80.5, 60.7, 38.7) (82.9, 67.8, 42.1) (82.8, 66.4, 40.9) (87.9, 77.4, 51.1)
TN3K (73.6, 37.8, 35.6) (81.5, 50.4, 47.9) (76.8, 34.0, 40.7) (82.4, 46.0, 46.5)
Average (80.5, 60.2, 43.9) (84.0, 67.0, 49.5) (83.1, 55.3, 47.6) (87.2, 71.4, 54.8)
Table 2: Ablation study key component. Performance is reported at image-level (AUROC, AP) and pixel-level (AUROC, AUPRO) on MVTec and VisA. Higher values indicate improved performance. Default configuration is highlighted.
(a) Prompting approaches effect.
Image-level Pixel-level
Prompts MVTec VisA MVTec VisA
Fixed (93.4, 96.1) (87.7, 90.9) (55.4, 24.6) (43.1, 11.5)
Learned (84.4, 92.9) (67.1, 72.9) (90.9, 84.0) (95.9, 88.2)
[…]
(b) Loss types effect in learning prompts.
Image-level Pixel-level
Loss MVTec VisA MVTec VisA
global (94.2, 97.2) (81.4, 84.4) (73.9, 53.1) (63.0, 36.7)
[…]
both (93.6, 97.0) (78.9, 82.6) (88.6, 77.4) (94.4, 84.9)
(c) Image-level strategies effect.
Image-level
St. Equation MVTec VisA
S1 $p_{a}(g_{o}^{i},G_{f})$ (93.5, 96.2) (82.9, 85.5)
S2 $p_{a}(g_{s}^{i},G_{f})$ (95.3, 97.6) (83.4, 85.8)
S3 $\max(S1,S2)$ (95.0, 97.2) (83.7, 86.3)
S4 $\text{mean}(S1,S2)$ (94.4, 96.5) (82.9, 85.8)
[…]
Implementation Details. We use the publicly available TIPS-L/14 HR [19] as our default backbone. It has 487.1M parameters, comparable to the CLIP-L/14@336 (428M) used in prior works. We use 8 learnable tokens for each of $T^{\mathrm{n}}$ and $T^{\mathrm{a}}$. For optimizer, we use Adam with a learning rate of 0.001 and betas=(0.5, 0.999), and train for 2 epochs with a batch size of 8. The random seed for all algorithms is set to 111 to ensure reproducibility. Input images to the vision tower are resized to 518×518 and undergo the same normalization during both training and inference. Experiments are conducted using PyTorch 2.5 on a NVIDIA A6000 GPU.
3.1 Quantitative Evaluations
Industrial Evaluation. We evaluate Tipsomaly against three CLIP-based baselines [34, 4, 5] on seven industrial datasets; results are summarized in Table 1. Averaged across datasets, Tipsomaly improves image-level AUROC, AP, and F1-max by +2.3%, +3.9%, and +1.1% over prior work, attaining the top result on most datasets. Notably, these gains are achieved without complex and computationally costly adaptation tricks. At the pixel level, Tipsomaly delivers even larger improvements in AUROC, AUPRO, and F1-max of +2.0%, +6.9%, and +1.5%, indicating stronger localization while maintaining a lightweight design.
Medical Evaluation. To assess cross-domain generalization beyond industrial data, we evaluate Tipsomaly on seven medical datasets; results are summarized in Table 1. At the pixel level, Tipsomaly improves AUROC, AUPRO, and F1-max by a remarkable margin of +3.2%, +4.4%, and +5.3% on average while remaining competitive at image-level, indicating broader understanding and strong zero-shot generalization across challenging medical benchmarks.
3.2 Qualitative Evaluations
The qualitative comparison in Fig. 2 shows that our method produces cleaner anomaly maps with sharply reduced false positives and more complete coverage of true anomalous regions. In contrast, competing approaches either over-highlight normal areas or miss critical defect regions.
3.3 Ablation Studies
In following ablations we evaluate TIPS and highlight effect of each introduced strategies to address the TIPS challenges.
Decoupled Prompting. Table 2(a) compares fixed textual prompts, learnable prompts, and a decoupled design. Fixed prompts favor image-level classification but collapse on segmentation, whereas learnable prompts excel at pixel-level localization yet underperform in classification. Decoupling resolves this inconsistency: Using fixed prompts for image-level scoring and learnable prompts for pixel-level maps, retaining strong image-level and pixel-level performance.
Learnable Prompts Objective. Table 2(b) examines global vs. local losses for learning prompts. A local loss alone yields the best pixel-level results. Adding a global classification loss improves image-level performance slightly but reduces pixel-level accuracy, and still trails the image-level results obtained with fixed prompts denoting an existing distributional gap between global and local embeddings. Hence, we rely on fixed prompts for image-level detection and optimize learnable prompts with the local loss for segmentation.
Image-level Strategies Table 2(c) assesses the discriminative ability of $g^{i}_{o}$ and $g^{i}_{s}$ for image-level anomaly detection, along with aggregation strategies (mean vs. max), with $p_{a}(g^{i}_{s},G_{f})$ yielding the best discrimination. Finally, augmenting the global score with the strongest local evidence, $\max(\hat{S}_{a})$, produces a net gain on the ten image-level datasets reported in Table 1, though this is not uniformly reflected by the two ablation sets.
Fig. 2: Qualitative results. Anomalies colored in original images. Tipsomaly yields cleaner localization with fewer false positives.
4 Conclusion
We investigated whether a spatially aware text–image backbone can deliver robust zero-shot anomaly detection without the intensive CLIP modifications common in prior work, demonstrating that TIPS can achieve this with a simple framework and fewer constraints. We showed a distributional gap between global and local visual embeddings in TIPS, which we bridged with decoupled prompting, resulting in the method Tipsomaly, which delivers state-of-the-art image- and pixel-level results across industrial and medical benchmarks, outperforming complex CLIP adaptations.
5 Acknowledgements
Mohammad Sabokrou’s work in this project was supported by JSPS KAKENHI Grant Number 24K20806.

### v2

@@
 When target-domain normal data are unavailable, zero-shot anomaly detection (ZSAD) leverages vision–language models (VLMs).
-However, CLIP’s coarse image–text alignment limits both localization and detection due to (i) spatial misalignment and (ii) weak sensitivity to fine-grained anomalies; prior work compensates with complex auxiliary modules yet largely overlooks the choice of backbone.
+However, CLIP’s coarse image–text alignment limits both localization and detection due to (i) spatial misalignment and (ii) weak sensitivity to fine-grained anomalies; prior works compensate with complex auxiliary modules yet largely overlook the choice of backbone.
 We revisit the backbone and use TIPS—a VLM trained with spatially aware objectives.
@@
 Although helpful, these adaptations cannot fully restore the spatial alignment and patch–text correspondence absent in CLIP inherently, and increase architectural complexity.
-Moreover, previous analyzes suggest that learnable adaptations can overfit train data, weakening the cross-domain generalization required for zero-shot anomaly detection [26].
+Moreover, previous analyses suggest that learnable adaptations can overfit train data, weakening the cross-domain generalization required for zero-shot anomaly detection [26].
 For example, AdaCLIP’s [4] joint tuning of text and vision encoders can improve image-level detection, but transfers poorly to pixel-level localization, underscoring the limits of CLIP-centric adaptations.
 In this paper, we revisit an underexplored lever for ZSAD—the backbone—and investigate TIPS’s [19] capabilities, a more spatially aware model for the task, showing it sidesteps several CLIP limitations and enables a simpler pipeline without complex adaptations.
-However, TIPS is trained with a generic vision–language objective, making direct utilization suboptimal for anomaly detection.
+However, TIPS is trained with a general vision–language objective, making direct utilization suboptimal for anomaly detection.
 We evaluated learnable prompts and fixed textual templates and found a consistent distributional gap between local and global features: when learnable prompts are trained with an additional global objective (normal vs.
@@
 $(\hat{y}(x),\,\hat{S}(x))\;=\;M_{\theta}(x,P).$
-thereafter the final Image- and pixel-level decisions are obtained by thresholding these scores.
+Thereafter, the final image- and pixel-level decisions are obtained by thresholding these scores.
 2.2 Feature Extraction
@@
 $P_{l}\;=\;\big\{\,[T^{\mathrm{n}},~\text{``object''}],\;[T^{\mathrm{a}},~\text{``damaged''},~\text{``object''}]\,\big\}.$
-Passing each element of $P_{l}$ through the text encoder $T(\cdot)$ yields the normal/abnormal localization prototypes $G_{l}=\{g^{n}_{l},\,g^{a}_{l}\}$ which used to score patch embeddings $Z^{\mathrm{M}}$ via similarity.
+Passing each element of $P_{l}$ through the text encoder $T(\cdot)$ yields the normal/abnormal localization prototypes $G_{l}=\{g^{n}_{l},\,g^{a}_{l}\}$ which are used to score patch embeddings $Z^{\mathrm{M}}$ via similarity.
 2.5 Training Objective & Inference
-Having obtained the textual embeddings specialized for detection $G_{f}$ and localization $G_{l}$ and visual embeddings $e\in\{g_{o}^{i},g_{s}^{i},Z_{j,k}\}$, where $Z_{j,k}$ denotes $(j_{th},k_{th})$ patch embedding of unflattened local feature $Z^{M}$.
-for input image $x$, we can now compute the likelihood of each visual embedding belonging to the anomalous class ($p_{a}$) by applying the Softmax function to similarity scores:
+Having obtained the textual embeddings specialized for detection $G_{f}$ and localization $G_{l}$, as well as the visual embeddings $e\in\{g_{o}^{i},g_{s}^{i},Z_{j,k}\}$, where $Z_{j,k}$ denotes the $(j,k)$-th patch embedding in the unflattened local feature map $Z^{M}$, we compute the probability that each visual embedding belongs to the anomalous class, denoted by $p_{a}$, by applying the Softmax function to the corresponding similarity scores:
 $p\!\left(e,G_{(\cdot)}\right)=\operatorname{softmax}\!\left(\frac{G_{(\cdot)}^{\top}e}{\tau}\right),\quad p_{a}=\bigl[p(e,G_{(\cdot)})\bigr]_{a}$ (1)
 where the temperature $\tau$ is approximately 0.0042 according to TIPS details [19]. We denote the probability of a visual embedding $e$ being abnormal, $p_{a}(e,G)$, as its anomaly score.
-Train.
-We learn the normal/abnormal patch-level visual distribution via learnable prompts $P_{l}$, while keeping the vision encoder $\Phi_{v}$ and text encoder $\Phi_{t}$ frozen and employing local loss function, Focal [18] and Dice [29] loss.
-For output feature map $Z^{M}$, we compute normal and anomaly maps $\hat{S}_{n}$ and $\hat{S}_{a}$ based on the equation 1, then apply bilinear upsampling $\text{Up}(\cdot)$ to match the anomaly mask $S\in\mathbb{R}^{H\times W}$:
+Training.
+We learn the normal/abnormal patch-level visual distribution via learnable prompts $P_{l}$, while keeping the vision encoder $\Phi_{v}$ and text encoder $\Phi_{t}$ frozen and employing local loss functions: Focal loss [18] and Dice loss [29].
+For output feature map $Z^{M}$, we compute normal and anomaly maps $\hat{S}_{n}$ and $\hat{S}_{a}$ based on Eq.
+(1), then apply bilinear upsampling $\text{Up}(\cdot)$ to match the anomaly mask $S\in\mathbb{R}^{H\times W}$:
 $L=\text{Focal}\left(\text{Up}([\hat{S}_{n},\hat{S}_{a}]),S\right)+\text{Dice}\left(\text{Up}([\hat{S}_{n},\hat{S}_{a}]),S\right).$
 Inference.
-For each input image $x$, after computing visual outputs including local features $Z^{M}$ and global embeddings $g_{o}^{i}$ and $g_{s}^{i}$ alongside detection and localization textual embeddings $G_{f}$ and $P_{l}$ in forward pass, low-resolution anomaly maps $\hat{S}_{a}$ is calculated then bilinearly upsampled and smoothed with Gaussian filter to obtain final anomaly map $\hat{S}$:
+For each input image $x$, after computing visual outputs including local features $Z^{M}$ and global embeddings $g_{o}^{i}$ and $g_{s}^{i}$ alongside detection and localization textual embeddings $G_{f}$ and $G_{l}$ in forward pass, low-resolution anomaly map $\hat{S}_{a}$ is calculated, then bilinearly upsampled and smoothed with Gaussian filter to obtain final anomaly map $\hat{S}$:
 $\hat{S}_{a,(j,k)}=p_{a}(Z^{M}_{j,k},G_{l}),\quad\hat{S}=Gaus(Up(\hat{S}_{a})).$
@@
 3 Experiments
-We evaluate on diverse 14 real-world anomaly detection datasets covering industrial defects and medical abnormalities, including MVTec-AD [2], VisA [35], DTD-Synthetic [1], SDD [30], BTAD [20], DAGM [33], and MPDD [14] for industrial and ISIC [10], CVC-ColonDB [31], CVC-ClinicDB [3], TN3K [8], BrainMRI [15], HeadCT [16], and BR35H [11] for medical domain.
+We evaluate on 14 diverse real-world anomaly detection datasets covering industrial defects and medical abnormalities, including MVTec-AD [2], VisA [35], DTD-Synthetic [1], SDD [30], BTAD [20], DAGM [33], and MPDD [14] for industrial and ISIC [10], CVC-ColonDB [31], CVC-ClinicDB [3], TN3K [8], BrainMRI [15], HeadCT [16], and BR35H [11] for medical domain.
 Following prior works [34, 4], we train on MVTec and test generalization on the others; for MVTec evaluation, we use VisA.
@@
 Average (80.5, 60.2, 43.9) (84.0, 67.0, 49.5) (83.1, 55.3, 47.6) (87.2, 71.4, 54.8)
-Table 2: Ablation study key component.
+Table 2: Ablation study of key components.
 Performance is reported at image-level (AUROC, AP) and pixel-level (AUROC, AUPRO) on MVTec and VisA.
@@
 Input images to the vision tower are resized to 518×518 and undergo the same normalization during both training and inference.
-Experiments are conducted using PyTorch 2.5 on a NVIDIA A6000 GPU.
+Experiments are conducted using PyTorch 2.5 on an NVIDIA A6000 GPU.
 3.1 Quantitative Evaluations
@@
 Fixed prompts favor image-level classification but collapse on segmentation, whereas learnable prompts excel at pixel-level localization yet underperform in classification.
-Decoupling resolves this inconsistency: Using fixed prompts for image-level scoring and learnable prompts for pixel-level maps, retaining strong image-level and pixel-level performance.
+Decoupling resolves this inconsistency: we use fixed prompts for image-level scoring and learnable prompts for pixel-level maps, retaining strong performance at both levels.
 Learnable Prompts Objective.
@@
 Hence, we rely on fixed prompts for image-level detection and optimize learnable prompts with the local loss for segmentation.
-Image-level Strategies Table 2(c) assesses the discriminative ability of $g^{i}_{o}$ and $g^{i}_{s}$ for image-level anomaly detection, along with aggregation strategies (mean vs.
+Image-level Strategies.
+Table 2(c) assesses the discriminative ability of $g^{i}_{o}$ and $g^{i}_{s}$ for image-level anomaly detection, along with aggregation strategies (mean vs.
 max), with $p_{a}(g^{i}_{s},G_{f})$ yielding the best discrimination.